# CHAIR-Classifier of Hallucination As Improver

### v1

@@ Abstract
This paper presents a supervised method for detecting hallucinations in large language models. By analyzing token scores (logitis) across layers of the LLaMA model (Touvron et al., ), we derive a small set, aiming to reduce overfitting, of features-including maximum, minimum, mean, standard deviation, and slope. We use logistic regression for classification and validate the model on the TruthfulQA and MMLU datasets. The results demonstrate significant performance gains, especially in zero-shot scenarios, highlighting the effectiveness and potential for generalization.
Machine Learning, ICML
1 Introduction
Hallucinations are a common issue in large language models (LLMs), limiting their reliability in critical applications, and a lot of research has been working on it (Huang et al., ) (Ji et al., ). Hallucinations usually manifest as factual inaccuracies, misleading information, or even entirely fabricated content in model-generated text. For instance, when asked, ”What did Einstein invent?” the model might hallucinate and incorrectly associate Einstein with inventions like the ”lightbulb” or ”telephone,” which are actually Edison’s inventions. This issue not only affects the credibility of the model, but also poses potential risks in fields like healthcare and law.
@@ 1 Introduction
1.1 Related Work
Recent studies have proposed various innovative strategies to reduce hallucinations. Factual nucleus sampling(Lee et al., ) is one approach that constrains word selection based on probabilities during decoding, increasing the likelihood of generating factual correct content. For example, in response to ”How many satellites does Earth have?” limiting the vocabulary selection can guide the model to generate a response like ”one” or ”the Moon,” improving factual accuracy. Contrastive decoding (Li et al., b) builds on this by contrasting different decoding paths to improve the reliability of the generated content. This approach focuses on tokens that initially receive low scores in the earlier layers but increase in score in later layers, suggesting that these tokens are syntactically plausible but semantically imprecise in the early stages, only becoming more accurate in later layers. For example, the model might initially lean toward generating ”Mars has plenty of water,” but later layers refine this to ”Mars has limited water resources.”
Another technique, DoLA(Chuang et al., ), examines how different attention-layers contribute to hallucination. By analyzing attention scores across multiple layers, DoLA highlights which layers to be ”contrasted” tend to produce more factual and reliable outputs. This method shows that hallucination can sometimes arise from attention imbalances across layers and suggests that integrating attention information from multiple layers, rather than just relying on the final layer, can improve model accuracy.
Inference-Time Intervention (Li et al., a) also addresses hallucinations by adjusting the model’s internal representations through vector shifting. For example, when answering ”What elements make up carbon dioxide?” this method can help the model accurately generate ”carbon and oxygen” rather than introducing unrelated elements. This approach aims to find a suitable adjustment direction during inference, helping align the internal representations with factual correctness.
In recent research, there has been increasing interest in leveraging the internal representations of LLMs. For example, (Skean et al., ) discusses how these internal layers can provide more informative features for downstream tasks compared to solely relying on the final layers. This insight highlights the potential benefits of exploring and utilizing information from internal representations, which aligns well with the objectives of our work in improving hallucination detection.
1.2 Motivation
Building on these approaches and analyzing the model’s internal behavior on TruthfulQA (Lin et al., ) shown in Figure 1, we propose a new perspective: intermediate layer information is not only useful for syntactic structure but also provides additional signals for semantic accuracy. Solely relying on the final layer might be insufficient for fully capturing hallucinations, while integrating intermediate layer information could enhance response accuracy.
@@ 1.2 Motivation
Using token-logit-traces, we performed supervised learning experiments on datasets such as TruthfulQA (Lin et al., ) and MMLU (Hendrycks et al., ) designing a logistic regression classifier that aggregates token scores across all layers in LLaMA (Touvron et al., ) or Mistral (Jiang et al., ) into a compact set of key features. This classifier is then used to detect hallucinations in model outputs. Experimental results show that our method significantly improves performance in these datasets, especially in multiple-choice tasks, where accuracy is notably enhanced. This approach validates the effectiveness of intermediate layer information in hallucination detection and provides a strong foundation for future applications in identifying and mitigating hallucinations across various use cases.
@@
2 Method
Our approach, named Classifier of Hallucination As Improver (CHAIR), leverages… to effectively detect hallucinations in model outputs.
CHAIR leverages representations generated by each layer of a model structure similar to LLaMA to identify hallucinations in language model outputs. LLaMA is a transformer-based model that applies self-attention (Vaswani et al., ) and feed-forward layers iteratively to build contextual representations of tokens. A crucial component in this setup is the use of an lm_head layer, which maps the hidden state from the last layer to a score, representing the model’s confidence in each token.
@@ 2 Method
2.1 Architecture of CHAIR
Figure 2 shows the CHAIR model structure for hallucination detection, adapted to handle multiple tokens at the feature extraction level. The input at the bottom includes an ”input embedding” consisting of a question and multiple ”choice tokens,” represented by a light gray rectangular box. Each choice token is highlighted in a different color (blue, red, green) to indicate separate tokens.
The model includes multiple layers, represented by yellow rectangular blocks labeled from ”layer-1” to ”layer-L.” Information flows sequentially through these layers, with arrows indicating the progression. After each layer processes the input, it passes the output to an $lm\_head$ module, shown as rounded rectangles positioned to the right of each layer block. Each $lm\_head$ module applies a transformation to the respective layer’s output, making it suitable for feature extraction.
After each $lm\_head$ transformation, the outputs are separated by token and directed to individual token-specific nodes represented by colored circles: red, green, and blue. Each colored circle corresponds to a specific token, and these token-specific feature nodes are connected to the $lm\_head$ outputs by colored arrows (red, green, blue) to maintain token-level distinction.
These token-specific feature nodes are aggregated in columns, each surrounded by a dashed box labeled ”Used to build features of the tokens.” This grouping indicates that the features from each token across all layers are combined to form a comprehensive feature set for each token.
Finally, the token-specific features are passed to the classifier, represented by an orange oval labeled ”Classifier.” The classifier takes the aggregated token-level features as input and uses them to determine whether the answer is hallucinatory. This result is indicated by an arrow leading to the text ”Is it a hallucination answer?” on the right.
In this enhanced design, the architecture maintains token-level separation throughout the layer-wise processing and feature extraction stages, allowing the model to capture nuanced patterns across different tokens and improve its accuracy in hallucination detection.
Figure 2: The diagram illustrates the CHAIR model architecture for hallucination detection, incorporating token-level features across multiple tokens. Yellow rectangles represent model layers layer-1 to layer-L, with $lm\_head$ modules shown as rounded rectangles. The circles red, green, blue denote feature nodes for individual tokens, with colors corresponding to different tokens. These features are aggregated through dashed lines and passed to the classifier orange oval, which determines if the answer is hallucinatory.
2.2 LLaMA-like Model Structure
@@
2.3 Token’s score history
@@ 6. Slope:
3 Experiment and Analysis
@@
3.1 Train and Test on Dataset
Table 1 presents a comparison of the performance of three approaches-baseline(Meta-Llama-3-8B-Instruct), DOLA, and our proposed method-across multiple metrics derived from the TruthfulQA and MMLU dataset. Each row corresponds to a specific metric, with columns showing the performance scores for each approach. Our method outperforms both Llama-3 and DOLA on the MC1 and MC3 metrics on TruthfulQA task, demonstrating a significant improvement in these areas. However, on the MC2 metrics, DOLA achieves relatively higher scores, though both DOLA and our method remain substantially better than the Llama-3.
Table 1: Comparative Performance of Mistral, DoLA, and Our Method across TruthfulQA and MMLU Metrics.
@@ 3.1 Train and Test on Dataset
3.2 Robustness of single dataset
@@
3.3 Generalization and Robustness cross multiple dataset
The Section 3.3 and Section 3.3 present the evaluation scores obtained when models are trained on one dataset (represented by rows) and evaluated on another dataset (represented by columns). Each cell $(i,j)$ indicates the change in performance when the model is trained on the dataset corresponding to row $i$ and tested on the dataset corresponding to column $j$. For example, the first column contains the change in MC1 scores, reflecting the variation in performance on TruthfulQA. In contrast, the second and third columns show the changes in average accuracy for the other datasets.
@@ 3.3 Generalization and Robustness cross multiple dataset
Table 2: Performance Gain over Meta-Llama-3-8B-Instruct Matrix: Training on Rows, Evaluating on Columns
@@
Table 3: Performance Gain over Mistral-7B-Instruct-v0.3 Matrix: Training on Rows, Evaluating on Columns

### v2

@@ Abstract
In this work, we introduce CHAIR (Classifier of Hallucination As ImproveR), a supervised framework for detecting hallucinations by analyzing internal logits from each layer of every token. Our method extracts a compact set of features—such as maximum, minimum, mean, standard deviation, and slope—from the token logits across all layers, enabling effective hallucination detection without overfitting. Experiments on TruthfulQA and MMLU datasets demonstrate that CHAIR significantly improves detection accuracy, particularly in zero-shot scenarios, showcasing its robustness and generalizability.
Beyond hallucination detection, CHAIR highlights the potential of using internal representations for designing advanced decoding strategies. By leveraging patterns in logits, we suggest that more sophisticated models and adaptive decoding methods could further reduce hallucinations and enhance text completion quality. CHAIR not only offers a practical solution for detecting hallucinations but also lays the groundwork for exploring richer representations in LLMs to improve their factuality and coherence.
Index Terms:
Large language models, Representation learning
I Introduction
Hallucinations are a common issue in large language models (LLMs), limiting their reliability in critical applications, and a lot of research has been working on it [1] [2]. Hallucinations usually manifest as factual inaccuracies, misleading information, or even entirely fabricated content in model-generated text. For instance, when asked, ”What did Einstein invent?” the model might hallucinate and incorrectly associate Einstein with inventions like the ”lightbulb” or ”telephone,” which are actually Edison’s inventions. This issue not only affects the credibility of the model, but also poses potential risks in fields like healthcare and law.
@@ I Introduction
I-A Related Work
Recent studies have proposed various innovative strategies to reduce hallucinations. Factual nucleus sampling[3] is one approach that constrains word selection based on probabilities during decoding, increasing the likelihood of generating factual correct content. For example, in response to ”How many satellites does Earth have?” limiting the vocabulary selection can guide the model to generate a response like ”one” or ”the Moon,” improving factual accuracy. Contrastive decoding [4] builds on this by contrasting different decoding paths to improve the reliability of the generated content. This approach focuses on tokens that initially receive low scores in the earlier layers but increase in score in later layers, suggesting that these tokens are syntactically plausible but semantically imprecise in the early stages, only becoming more accurate in later layers. For example, the model might initially lean toward generating ”Mars has plenty of water,” but later layers refine this to ”Mars has limited water resources.”
Another technique, DoLA[5], examines how different attention-layers contribute to hallucination. By analyzing attention scores across multiple layers, DoLA highlights which layers to be ”contrasted” tend to produce more factual and reliable outputs. This method shows that hallucination can sometimes arise from attention imbalances across layers and suggests that integrating attention information from multiple layers, rather than just relying on the final layer, can improve model accuracy.
Inference-Time Intervention [6] also addresses hallucinations by adjusting the model’s internal representations through vector shifting. For example, when answering ”What elements make up carbon dioxide?” this method can help the model accurately generate ”carbon and oxygen” rather than introducing unrelated elements. This approach aims to find a suitable adjustment direction during inference, helping align the internal representations with factual correctness.
In recent research, there has been increasing interest in leveraging the internal representations of LLMs. For example, [7] discusses how these internal layers can provide more informative features for downstream tasks compared to solely relying on the final layers. This insight highlights the potential benefits of exploring and utilizing information from internal representations, which aligns well with the objectives of our work in improving hallucination detection.
I-B Motivation
Building on these approaches and analyzing the model’s internal behavior on TruthfulQA [8] shown in Figure 1, we propose a new perspective: intermediate layer information is not only useful for syntactic structure but also provides additional signals for semantic accuracy. Solely relying on the final layer might be insufficient for fully capturing hallucinations, while integrating intermediate layer information could enhance response accuracy.
@@ I-B Motivation
Using token-logit-traces, we performed supervised learning experiments on datasets such as TruthfulQA [8] and MMLU [9] designing a logistic regression classifier that aggregates token scores across all layers in LLaMA [10] or Mistral [11] into a compact set of key features. This classifier is then used to detect hallucinations in model outputs. Experimental results show that our method significantly improves performance in these datasets, especially in multiple-choice tasks, where accuracy is notably enhanced. This approach validates the effectiveness of intermediate layer information in hallucination detection and provides a strong foundation for future applications in identifying and mitigating hallucinations across various use cases.
@@
II Method
Our approach, named Classifier of Hallucination As Improver (CHAIR), leverages representations generated by each layer of a model structure similar to LLaMA to identify hallucinations in language model outputs. LLaMA is a transformer-based model that applies self-attention [12] and feed-forward layers iteratively to build contextual representations of tokens. A crucial component in this setup is the use of an lm_head layer, which maps the hidden state from the last layer to a score, representing the model’s confidence in each token.
@@ II Method
II-A Architecture of CHAIR
Before we delve into the CHAIR architecture, let’s first revisit the structure of a LLaMA-like model. At its core, input tokens are converted into embeddings and passed through multiple layers of transformer blocks. Each block consists of layer normalization, a multi-head attention mechanism, and a feed-forward layer that processes each token independently. Layer normalization is applied before both the multi-head attention and feed-forward layers. Residual connections are incorporated to ensure stable training. The output from the final transformer layer is passed to a projection layer, referred to as the $lm\_head$.
Figure 2 illustrates the CHAIR model architecture, which is designed to handle multiple tokens at the feature extraction level. At the bottom, the input consists of an ”input embedding” comprising a question represented by a light gray rectangular box and multiple ”choice/answer tokens”. Each choice token is highlighted in a different color (e.g., blue and green) to signify distinct tokens.
The model contains multiple transformer blocks, identical to those in the LLaMA architecture, depicted as gray boxes labeled ”block-1” to ”block-N.” The output of each block is passed to an $lm\_head$ module, which applies a transformation specific to the corresponding block’s output.
After the $lm\_head$ transformation, the resulting outputs are treated as features for the corresponding token at that layer or block. Each colored feature corresponds to a specific token.
These token-specific features are aggregated and grouped, as indicated by dashed boxes. This grouping signifies that the features from all layers for each token are combined to construct a comprehensive feature set for that token.
Finally, the aggregated token-specific features are passed to a classifier. The classifier processes these features to determine whether the generated answer is hallucinatory.
Figure 2: This illustrates the structure of CHAIR. The blue and green represent the tokens for choices/answers, while the gray boxes denote the transformer layers in the LLaMA model. The purple rectangles correspond to the references of the $lm\_head$ layer, and the yellow rectangle represents the feature extraction module. The classifier processes the inputs to determine whether the output sequence represents a hallucination.
II-B LLaMA-like Model Structure
@@
II-C LM Head and History of Token’s Score
@@ 6. Slope:
II-E Supervised Learning
In this section, we describe the process of training the classifier. Constructing the training and testing datasets is relatively straightforward: for a given question, the expected answers are labeled as 1, while all other answers are labeled as 0.
For instance, as shown in Figure 1, the green answers are labeled as 1, whereas the red answers are labeled as 0.
The challenging part lies in how to sample from the dataset. We explored several approaches:
1. Dataset-based sampling: Following the method in [6], we sampled directly from the dataset used for testing. This approach resembles an n-shot learning paradigm: we randomly sample $n$ examples to create a training set, train the classifier on this set, and then evaluate its performance on the test set. This method is compared against other n-shot approaches.
2. Impact of $n$: To determine the effect of the number of samples $n$ on classifier performance, we conducted multiple experiments using the same $n$. The results are analyzed to identify the minimum $n$ required for robust classifier performance.
3. Cross-dataset generalization: We hypothesize that the logits from internal layers of the LLM reflect intrinsic characteristics of the model, which should generalize across datasets. Under this assumption, the logits themselves can reveal whether an output is hallucinatory. Once the classifier learns this pattern from one dataset, it should perform well on others, validating the generalizability of the approach.
These experiments collectively evaluate the effectiveness and robustness of the classifier in detecting hallucinations in LLM outputs.
III Experiment and Analysis
@@
III-A Supervised Learning on Dataset
Table I presents a comparison of the performance of three approaches-baseline(Meta-Llama-3-8B-Instruct), DOLA, and our proposed method-across multiple metrics derived from the TruthfulQA and MMLU dataset. Each row corresponds to a specific metric, with columns showing the performance scores for each approach. Our method outperforms both Llama-3 and DOLA on the MC1 and MC3 metrics on TruthfulQA task, demonstrating a improvement in these areas. However, on the MC2 metrics, DOLA achieves relatively higher scores, though both DOLA and our method remain substantially better than the Llama-3.
TABLE I: Comparative Performance of Mistral, DoLA, and Our Method across TruthfulQA and MMLU Metrics.
@@ III-A Supervised Learning on Dataset
III-B Robustness on single dataset
@@
III-C Generalization and Robustness cross multiple dataset
The Section III-C and Section III-C present the evaluation scores obtained when models are trained on one dataset (represented by rows) and evaluated on another dataset (represented by columns). Each cell $(i,j)$ indicates the change in performance when the model is trained on the dataset corresponding to row $i$ and tested on the dataset corresponding to column $j$. For example, the first column contains the change in MC1 scores, reflecting the variation in performance on TruthfulQA. In contrast, the second and third columns show the changes in average accuracy for the other datasets.
@@ III-C Generalization and Robustness cross multiple dataset
TABLE II: Performance Gain over Meta-Llama-3-8B-Instruct Matrix: Training on Rows, Evaluating on Columns
@@
TABLE III: Performance Gain over Mistral-7B-Instruct-v0.3 Matrix: Training on Rows, Evaluating on Columns
@@
IV Conclusion
In this work, we presented CHAIR, a framework for detecting hallucinations in large language models (LLMs) by leveraging internal logits from each layer of every token. By using a lightweight supervised learning classifier, CHAIR demonstrated significant improvements in hallucination detection on datasets like TruthfulQA and MMLU, particularly excelling in zero-shot scenarios.
While the current approach has proven effective, there remains considerable potential for further advancements. The internal logits extracted from each layer offer rich, hierarchical representations that could be better exploited with more sophisticated models, such as transformer-based classifiers. These models might capture more complex interactions between layers and tokens, potentially improving classification accuracy and robustness across diverse datasets.
Moreover, our findings suggest that understanding the internal dynamics of LLMs could pave the way for developing better decoding strategies. By incorporating insights from logits patterns into decoding algorithms, we can guide the model toward generating more factual and coherent completions, reducing the likelihood of hallucinations during inference. Techniques such as adaptive decoding or contrastive reranking could integrate seamlessly with CHAIR to further enhance its utility in real-world applications.
By extending CHAIR with advanced models and decoding techniques, we can continue to improve the reliability and trustworthiness of language model outputs, enabling their broader adoption in critical domains.